Find the area bounded by the curve y = |x−1|, the x-axis and the lines x = −7 and x = 11
1
Expert's answer
2020-06-08T18:02:05-0400



"\text{As shown in the figure},\\
\text{ the area is divided to parts}\\

A=\int\limits_{-7}^1(-x+1-0)dx+\int\limits_1^{11}(x-1-0)dx\\

=\int\limits_{-7}^1(-x+1)dx+\int\limits_1^{11}(x-1)dx\\

=(-x^2/2+x)|_{-7}^1+(x^2/2-x)|_1^{11}\\

=-0.5+1+49/2+7\\
+121/2-11-0.5+1\\

=82"